# PPTA Collaboration

### v1

@@ VII.1 Time series of the PA residuals
Table 2: Performance analysis of different noise models for the 22 PPPA MSPs, using the data after ionospheric subtraction. The four “Posterior” columns display the posterior distributions of the white-noise and red-noise parameters in the full noise model, where the unconstrained parameters are marked as “…”. The last four columns show the Bayes factors for model comparison. Here, $\psi^{(0)}$, $\psi^{(0,1)}$, and $\psi^{(0,1,2)}$ denote the noise models with the minimal white noise of $\sigma_{p,n}$ for $\Delta{\rm PA}^{\rm obs}$ and the polynomial deterministic noise up to the $0^{\rm th}$, $1^{\rm st}$, and $2^{\rm nd}$ orders, respectively. The notations of “$+w$” and “$+r$” indicate the incorporation of additional white-noise and red-noise components in the noise model, respectively.
We then perform a Bayesian analysis with $\mathcal{L}_{m}$ to examine the performance of the noise models, by fitting them to the data of 22 PPPA MSPs after ionospheric subtraction using ionFR. The results are summarized in Table 2. To compare the performance of these models, we compute Bayes factors $\textrm{BF}_{j}^{i}=Z_{i}/Z_{j}$, where $Z_{i}$ represents the evidence of the $i$-th model and is derived using the Markov Chain Monte Carlo (MCMC) chains generated by the parallel tempering MCMC package PTMCMCSampler Ellis and van Haasteren (2017) with thermodynamical integral Goggans and Chi (2004); Calderhead and Girolami (2009).
@@
Evidently, the additional noise components in the full noise model in Eq. (3) improve the data fit for the majority of the 22 pulsars, compared to the minimal noise model which includes only radiometer noise and a constant offset. The improvement is especially significant for pulsars such as J0437-4715, J1713+0747, and J1909-3744. The relatively narrow posterior distributions for the random noise parameters in the full noise model also reflect the reasonableness of this model. The contribution of each new component varies for different pulsars. For most pulsars, the additional white noise component leads to more noticeable improvements, suggesting the possible presence of ionospheric Faraday rotation, caused by imperfect modeling, on timescales shorter than the observation cadence. Regarding red noise, five of the seven pulsars that show positive evidence have relatively low $\widetilde{\rm DM}$ and RM values ($|\rm{RM}|<10$ rad m-2), indicating that the observed red noise might receive certain contributions from the variation of ionospheric effects over years. This conjecture is also consistent with the observation (see Table 1) that the measurement errors of PA residuals for these MSPs tend to be small, which may enhance the chance of recognizing the red noise contributions from the ionospheric effects in the PA residuals. The two exceptional pulsars, namely J1545--4550 and J1643--1224, have larger $\widetilde{\rm DM}$ values and might be subject to stronger ISM influences. Moreover, several pulsars exhibit strong evidence of linear and quadratic variations over long periods, as additional deterministic noise contributions. However, more precise RM measurements are needed for better understanding of their nature.
@@
We will utilize this advanced noise model for the data analysis generally. However, including red noise in the combined sensitivity analysis can significantly increases the numerical cost. The larger $k_{{\rm max},p}$ is, the longer it takes to calculate the inverse of the covariance matrix. So, we choose $k_{{\rm max},p}=30$ to accurately determine the red noise parameters $\Gamma_{p}$ and $S_{p}^{r}$ in the single pulsar analysis, and include the red noise component only for the pulsars with $\ln{\rm BF}^{\psi^{(0,1,2)}+w+r}_{\psi^{(0,1,2)}+w}>2.3$, namely J0437-4715, J1545-4550, J1713+0747 and J1909-3744, in the PPPA analysis. In addition, in the combined analysis, we choose $k_{{\rm max},p}$ according to the best-fit $\Gamma_{p}$ value from the single pulsar noise analysis. Specifically, we choose $k_{{\rm max},p}=5$ if $\Gamma_{p}<-3$, and $k_{{\rm max},p}=30$ if $\Gamma_{p}\geq-3$ to ensure that Eq. (24) converges to Eq. (22) sufficiently.
@@
Let us consider the covaraint matrix for the red noise first, and decompose it as a matrix multiplication:
@@
Figure 8 displays the 95% upper limits on $S_{a}$ as a function of $m_{a}$ for each of the 22 PPPA MSPs, supplementing Fig. 1 in the main text. The difference between the black and colored solid lines reflects the impacts of individual noise components on the $S_{a}$ limits. For each component, such impacts may vary a lot among different pulsars. Specifically, the ionospheric subtraction using ionFR effectively removes sharp peaks around $\sim 1\,$yr for J1600-3053, J1643-1224, J1730-2304, J1744-1134, J1824-2452 and J1939+2134, and reduces excess low-frequency noises for J0437-4715, J1024-0719, J1600-3053 and J1744-1134. Incorporating the additional contributions of white noise described by the parameters EFp and EQp in Eq. (20) strengthens the limits significantly from J1713+0747, J1730-2304, J1744-1134, J1939+2134, etc. Incorporating the red noise component notably improves the limits from J0437-4715 and J1909-3744 at both intermediate and low frequencies. To account for long-term deterministic effects, we conservatively apply a second-order polynomial model to all pulsars, although strong evidence of linear and quadratic temporal variations has been observed for some of them only. This strategy results in less stringent limits at frequencies below $1/T_{\rm obs}$, a small portion of the frequency range of interest. Overall, for the most sensitive pulsars, such as J0437-4715, J1744-1134, etc., these outcomes imply that using the minimal noise model may result in an underestimation of sensitivity by several times and even over an order of magnitude.
@@
Figure 9: Mock limits on the ALDM Chern-Simons coupling $g_{a\gamma\gamma}$ and the mock performance of the Bayes factors $\ln{\rm BF}^{\rm Full}_{\rm Null}$ and $\ln{\rm BF}^{\rm Full}_{\rm Auto}$ as a function of $m_{a}$, in three different signal scenarios. In the top panel, the mock data are generated with white and red noises only, based on parameter values derived from fitting the real data. In the middle and bottom panels, a stochastic ALDM signal is injected into these random noises, with the $\{m_{a},g_{a\gamma\gamma}\}$ values denoted by yellow star in each panel. The vertical solid line denotes the PPPA observation period $T_{\rm obs}\approx 18~{}{\rm yrs}$, and the dotted lines for other reference time scales, including solar cycle, one year and one day.
@@
Figure 9 illustrates the PPA limits on $g_{a\gamma\gamma}$ and the $\ln{\rm BF}^{\rm Full}_{\rm Null}$ and $\ln{\rm BF}^{\rm Full}_{\rm Auto}$ responses to the ALDM signals, using mock data for several typical examples in three different signal scenarios. For the mock data containing the noises only (top panel), both Bayes factors fluctuate around zero, indicating no clear evidence for a signal with either full correlation or auto-correlation only. However, with the ALDM signals injected, these two Bayes factors demonstrate some dramatical deviation from the baseline of $\ln{\rm BF}\sim 0$. If the ALDM mass is larger than $T_{\rm obs}$, allowing the signal to be well-resolved, a sharp peak appears in both Bayes factors at the true mass value (middle panel). If the ALDM mass is comparable to or smaller than $T_{\rm obs}$, the signal cannot be fully resolved, resulting in an upward deformation in both Bayes factors below the true mass value (bottom panel). As $\ln{\rm BF}^{\rm Full}_{\rm Null}$ measures the deviation from null signal scenario and $\ln{\rm BF}^{\rm Full}_{\rm Auto}$ examines the role of cross correlation in recognizing the signal, these features can serve as key signatures for the ALDM detection.
@@
Figure 10: Mock distributions of the Bayes factor $\ln{\rm BF}^{\rm Full}_{\rm Auto}$ in three different signal scenarios. In the top panel, the mock data are generated with white and red noises only, based on parameter values derived from fitting the real data. In the middle and bottom panels, a stochastic ALDM signal is injected into these random noises, with $m_{a}=10^{-21}\,$eV and $S_{a}=0.1\,$deg and 0.30.30.3\,0.3deg, respectively. 600 random realizations are analyzed for each case, all at $m_{a}=10^{-21}\,$eV.
@@
To statistically support this argument, we particularly display in Fig. 10 the mock distributions of the Bayes factor $\ln{\rm BF}^{\rm Full}_{\rm Auto}$ for three different signal scenarios, based on totally 1800 random realizations of both noise and signal. For the case with random noises only, the Bayes factor distribution, which is calculated at $m_{a}=10^{-21}\,$eV for the convenience of discussions, is centered around zero. In the cases with injected signals, we consider an ALDM mass larger than $T_{\rm obs}$, namely $m_{a}=10^{-21}\,$eV, for demonstration. The Bayes factor $\ln{\rm BF}^{\rm Full}_{\rm Auto}$ is then derived at this mass value to plot the distributions. As the signal strength, i.e., the $S_{a}$ value, increases, the distribution shifts to the positive side, with increased median and variance. This outcome further highlights the role of $\log{\rm BF}^{\rm Full}_{\rm Auto}$ or the cross-correlation in recognizing the nature of potential signals in the data.
@@
We present the PPPA limits on the ALDM Chern-Simons coupling in Fig. 2 for the full-correlation and auto-correlation-only signal models. Both sets of limits display sharp peaks on the timescales of one day or shorter and at $m_{a}\approx 10^{-22}\,$eV. To acquire insights regarding the nature of these peaks, we demonstrate both Bayes factors, namely $\ln{\rm BF}^{\rm Full}_{\rm Null}$ and $\ln{\rm BF}^{\rm Full}_{\rm Auto}$, as functions of $m_{a}$ in the same figure. Notably, while the peaks also appear in the Bayes factor $\ln{\rm BF}^{\rm Full}_{\rm Null}$ on the similar timescales, $\ln{\rm BF}^{\rm Full}_{\rm Auto}$ remains near zero across the entire mass range. These small values of $\ln{\rm BF}^{\rm Full}_{\rm Auto}$ are consistent with the conservative noise prediction within a single mass bin, as shown in the upper panel of Fig. 10. This suggests that no evidence for the ALDM-induced cross-correlation or signal has been found in the analyzed data, and the sharp peaks observed in the limits and also the $\log{\rm BF}^{\rm Full}_{\rm Null}$ curve are more likely caused by imperfect noise modeling for individual pulsars.

### v2

@@ VII.1 Time series of the PA residuals
Table 2: Performance analysis of different noise models for the 22 PPPA MSPs, using the data after ionospheric subtraction. The four “Posterior” columns display the posterior distributions of the white-noise and red-noise parameters in the full noise model, where the unconstrained parameters are marked as “…”. The last four columns show the Bayes factors for model comparison. Here, $\psi^{(0)}$, $\psi^{(0,1)}$, and $\psi^{(0,1,2)}$ denote the noise models with the minimal white noise of $\sigma_{p,n}$ for $\Delta{\rm PA}^{\rm obs}$ and the polynomial noise term up to the $0^{\rm th}$, $1^{\rm st}$, and $2^{\rm nd}$ orders, respectively. The notations of “+w+w” and “+r+r” indicate the incorporation of additional white-noise and red-noise components in the noise model, respectively.
[…]
@@
Evidently, the additional components (or at least one of them) in the full noise model (see Eq. (3)) improve the data fit for most of the 22 pulsars, likely excluding PSRs J1045–4509, J1643–1224, and J1824–2452, when compared to the minimal noise model. The improvement is especially significant for pulsars such as J0437–4715, J1713+0747, and J1909–3744. The relatively narrow posterior distributions for the random noise parameters in the full noise model also reflect the reasonableness of this model. The contribution of each new component varies for different pulsars. For most pulsars, the additional white noise component leads to noticeable improvements, suggesting the possible presence of ionospheric Faraday rotation, caused by imperfect modeling, on timescales shorter than the observation cadence. Several pulsars show evidence or strong hints of red noise, including J0437–4715, J1022–1001, J1545–4550, and J1909–3744. Additionally, despite the reduction of their annual modulation in PA residuals through ionospheric subtraction, J1713+0747 and J1744–1134 also exhibit this characteristic. The ISM variations may contribute significantly to this feature, particularly for PSR J1545-4550 which has the highest-$\widetilde{\rm DM}$ value among these six pulsars (see Table 1). Since we primarily consider the long-term effects by setting $f_{H}\leq 30f_{L}$ with $f_{L}=1/T_{{\rm obs},p}$ for the red-noise model given in Eq. (22), the timescales for the observed red noise might be years-long. This conjecture gets supported by our examination of the $\Delta{\rm PA}^{\rm obs}$ structure functions for these pulsars, using a method from Kumamoto et al. (2021). There is no power-law rise indicated in these functions. However, we did observe features over timescales of several years. These timescales exceed typical timescales of diffractive scintillation in the ISM Keith et al. (2013), and fall within the range associated with ISM refractive scintillation Kumamoto et al. (2021) and the evolution of pulsar RM Keith et al. (2024). Therefore, the observed red noise may arise from ISM structure and turbulence or fluctuation on these scales, as well as variations in the interstellar magnetic field along the line of sight to these pulsars. Notably, the relatively low $\widetilde{\rm DM}$ and RM values ($|\rm{RM}|<10$ rad m-2) for the pulsars such as J0437–4715 and J1909–3744 may suggest that the observed red noise also receives contributions from the unmodeled ionospheric effects over the years. Finally, several pulsars exhibit strong evidence of linear and quadratic variations over long periods, as additional noise contributions. Similar trends have been observed also in the DM and RM evolution of pulsars Keith et al. (2024); Lam et al. (2016).These outcomes reasonably justify our noise modeling.
@@
We will utilize this advanced noise model for the data analysis generally. However, including red noise in the combined sensitivity analysis can significantly increases the numerical cost. The larger $k_{{\rm max},p}$ is, the longer it takes to calculate the inverse of the covariance matrix. So, we choose $k_{{\rm max},p}=30$ to accurately determine the red noise parameters $\Gamma_{p}$ and $S_{p}^{r}$ in the single pulsar analysis, and include the red noise component only for the pulsars with $\ln{\rm BF}^{\psi^{(0,1,2)}+w+r}_{\psi^{(0,1,2)}+w}>2.3$, namely J0437–4715, J1545–4550, J1713+0747 and J1909–3744, in the PPPA analysis. In addition, in the combined analysis, we choose $k_{{\rm max},p}$ according to the best-fit $\Gamma_{p}$ value from the single pulsar noise analysis. Specifically, we choose $k_{{\rm max},p}=5$ if $\Gamma_{p}<-3$, and $k_{{\rm max},p}=30$ if $\Gamma_{p}\geq-3$ to ensure that Eq. (24) converges to Eq. (22) sufficiently.
Stochastic spacetime dependence of the ALDM field
[…]
Figure 9: Spatial dependence of the normalized ALDM field $\bar{a}(x)=a(\mathbf{x},t=0)m_{a}/\sqrt{\rho_{0}}$ in the direction of $\mathbf{x}$ ($x=|\mathbf{x}|$), for a specific realization of the random parameters $\{\alpha_{\mathbf{v}},\phi_{\mathbf{v}}\}$.
However, the pulsar distances to another pulsar or to the Earth can exceed the coherence length $l_{c}$. To preserve spatial information of the ALDM waves, we retain the leading term of the $\mathbf{k}\cdot\mathbf{x}$ non-relativistic expansion, namely $m_{a}\mathbf{v}\cdot\mathbf{x}$, in Eq. (1). This term accounts for the spatial coherence within the range of coherence length as well as its loss over distances larger than $l_{c}$. For illustration, we show in Fig. 9 the spatial variation of the normalized ALDM field, i.e., $\bar{a}(\mathbf{x},t)=a(\mathbf{x},t)m_{a}/\sqrt{\rho_{0}}$, for one random realization of $\{\alpha_{\mathbf{v}},\phi_{\mathbf{v}}\}$. Given the isotropy of the velocity distribution $f(\mathbf{v})$, we plot this variation in the direction of $\mathbf{x}$ ($x=|\mathbf{x}|$). The field variation approximately follows a coherent trigonometric function when $x\ll l_{c}$, and becomes stochastic for $x\gg l_{c}$. The ALDM with $m_{a}\sim 10^{-22}\,$eV has a coherence length of $l_{c}\sim\mathcal{O}(100)\,$pc. This number scales with $l_{c}\propto 1/m_{a}$ for other mass values. Note that this figure has been normalized and is applicable to the entire mass of ultralight ALDM.
@@
Let us consider the covariance matrix for the red noise first, and decompose it as a matrix multiplication:
@@
Fig. 10 displays the 95% upper limits on $S_{a}$ as a function of $m_{a}$ for each of the 22 PPPA MSPs, supplementing Fig. 1 in the main text. The difference between the black and colored solid lines reflects the impacts of individual noise components on the $S_{a}$ limits. For each component, such impacts may vary a lot among different pulsars. Specifically, the ionospheric subtraction using ionFR effectively removes sharp peaks around $\sim 1\,$yr for J1600–3053, J1643–1224, J1730–2304, J1744–1134, J1824–2452 and J1939+2134, and reduces excess low-frequency noises for J0437–4715, J1024–0719, J1600–3053 and J1744–1134. Incorporating the additional contributions of white noise described by the parameters EFp and EQp in Eq. (20) strengthens the limits significantly from J1713+0747, J1730–2304, J1744–1134, J1939+2134, etc. Incorporating the red noise component notably improves the limits from J0437–4715 and J1909–3744 at both intermediate and low frequencies. To account for long-term deterministic effects, we conservatively apply a second-order polynomial model to all pulsars, although strong evidence of linear and quadratic temporal variations has been observed for some of them only. This strategy results in less stringent limits at frequencies below $1/T_{\rm obs}$, a small portion of the frequency range of interest. Overall, for the most sensitive pulsars, such as J0437–4715, J1744–1134, etc., these outcomes imply that using the minimal noise model may result in an underestimation of sensitivity by several times and even over an order of magnitude.
@@
To demonstrate this point, we perform a mock study by applying the PPA analysis to the noise-only model and the noise plus ALDM signal models, respectively. The former case includes only white and red noises, with the model parameter values derived from fitting real data. The latter case incorporates the injected signals additionally. Here, we consider two ALDM benchmark scenarios: $\{m_{a},g_{a\gamma\gamma}\}$ = $\{10^{-21}\,{\rm eV},10^{-12}{\rm GeV}^{-1}\}$ and $\{10^{-23}\,{\rm eV},10^{-14}{\rm GeV}^{-1}\}$, with the corresponding characteristic signal strength $S_{a}$ being approximately 0.10.1\,deg for both.
[…]
@@
Figure 12: Mock PPA limits on the ALDM Chern-Simons coupling $g_{a\gamma\gamma}$ and mock performance of the Bayes factors $\ln{\rm BF}^{\rm Full}_{\rm Null}$ and $\ln{\rm BF}^{\rm Full}_{\rm Auto}$, as a function of the ALDM mass $m_{a}$. The top panel displays a sample of random noise only, while the bottom panels show two samples with the injected ALDM signals from the benchmark scenarios defined in Fig. 11. Here the yellow stars indicate the corresponding $\{m_{a},g_{a\gamma\gamma}\}$ values for these two scenarios. The vertical solid line denotes the PPPA observation period $T_{\rm obs}\approx 18~{\rm yrs}$, and the dotted lines for other reference timescales, including solar cycle, one year and one day.
@@
We first illustrate in Fig. 12 the PPA limits on $g_{a\gamma\gamma}$ and the $\ln{\rm BF}^{\rm Full}_{\rm Null}$ and $\ln{\rm BF}^{\rm Full}_{\rm Auto}$ responses to the ALDM signals for the considered samples of mock data. For the mock data containing the random noises only (top panel), both Bayes factors fluctuate around zero, indicating no clear evidence for a signal with either full correlation or auto-correlation only. However, with the ALDM signals injected, these two Bayes factors demonstrate some dramatical deviation from the baseline of $\ln{\rm BF}\sim 0$, exceeding the detection threshold of 2.3 for strong evidence. If the ALDM mass is larger than $1/T_{\rm obs}$, allowing the signal to be well-resolved, a sharp peak appears in both Bayes factors at the true mass value (bottom-left panel). Building on this, the strength of injected signals $S_{a}$ indeed can be recovered with a high precision, as illustrated by the mock posterior distributions in Fig. 13, which exhibit minimal degeneracy with noise parameters. If the ALDM mass is comparable to or smaller than $1/T_{\rm obs}$, the signal cannot be fully resolved, resulting in an upward deformation in both Bayes factors below the true mass value (bottom-right panel). As $\ln{\rm BF}^{\rm Full}_{\rm Null}$ measures the deviation from the null signal scenario and $\ln{\rm BF}^{\rm Full}_{\rm Auto}$ examines the role of cross correlation in recognizing the signal, these features can serve as key signatures for the ALDM detection. These discussions and observations clearly indicate that the ALDM signals can be efficiently detected through the PPA analysis, despite the incorporation of additional components into the full noise model. We emphasize that this outcome is to be expected, given the unique correlation pattern of the ALDM signals which can differentiate them from the noise.
Figure 13: Mock posterior distributions for the random noise parameters of the most sensitive pulsar, i.e., J0437–4715, and the characteristic signal strength $S_{a}$ from the PPA analysis at $m_{a}=10^{-21}\,$eV. The red contours, corresponding to the bottom-left panel of Fig. 12, represent the ALDM benchmark scenario with $\{m_{a},g_{a\gamma\gamma}\}$ = $\{10^{-21}\,{\rm eV},10^{-12}{\rm GeV}^{-1}\}$ (where $S_{a}\approx 0.1\,$deg), while the blue contours illustrate the results for the same sample but with no ALDM signals injected. The vertical or horizontal pink lines indicate the $S_{a}$ value for the injected ALDM signals.
@@
Figure 14: Mock distributions of the Bayes factor $\ln{\rm BF}^{\rm Full}_{\rm Auto}$ in three different signal scenarios. In the top panel, the mock data are generated with white and red noises only, based on parameter values derived from fitting the real data. In the middle and bottom panels, a stochastic ALDM signal is injected into these random noises, with $m_{a}=10^{-21}\,$eV and $S_{a}=0.1\,$deg and 0.30.3\,deg, respectively. 600 random realizations are analyzed for each case, all at $m_{a}=10^{-21}\,$eV.
@@
So far, only one specific realization for random noises and stochastic ALDM signals has been applied for the analysis of mock data. To statistically support this argument, we particularly display in Fig. 14 the mock distributions of the Bayes factor $\ln{\rm BF}^{\rm Full}_{\rm Auto}$ for three different signal scenarios, based on totally 1800 random realizations of both noises and signals. For the case with random noises only, the Bayes factor distribution, which is calculated at $m_{a}=10^{-21}\,$eV for the convenience of discussions, is centered around zero. In the cases with injected signals, we consider an ALDM mass larger than $1/T_{\rm obs}$, namely $m_{a}=10^{-21}\,$eV, for demonstration. The Bayes factor $\ln{\rm BF}^{\rm Full}_{\rm Auto}$ is then derived at this mass value to plot the distributions. As the signal strength, i.e., the $S_{a}$ value, increases, the distribution shifts to the positive side, with increased median and variance. This outcome further highlights the role of $\log{\rm BF}^{\rm Full}_{\rm Auto}$ or the pulsar cross-correlation in recognizing the real signals in the data.
@@
We present the PPPA limits on the ALDM Chern-Simons coupling in Fig. 2 for the full-correlation and auto-correlation-only signal models. Both sets of limits display sharp peaks on the timescales of one day or shorter and at $m_{a}\approx 10^{-22}\,$eV. To acquire insights regarding the nature of these peaks, we demonstrate both Bayes factors, namely $\ln{\rm BF}^{\rm Full}_{\rm Null}$ and $\ln{\rm BF}^{\rm Full}_{\rm Auto}$, as functions of $m_{a}$ in the same figure. Notably, while the peaks also appear in the $\ln{\rm BF}^{\rm Full}_{\rm Null}$ curve on the similar timescales, $\ln{\rm BF}^{\rm Full}_{\rm Auto}$ remains near zero across the entire mass range (similar to the $\ln{\rm BF}^{\rm Full}_{\rm Auto}$ shown in the upper panel of Fig. 12 and also consistent with its distribution shown in the upper panel of Fig. 14). This suggests that no evidence for the ALDM-induced cross-correlation or signal has been found in the analyzed data, and the sharp peaks observed in the limits and also the $\log{\rm BF}^{\rm Full}_{\rm Null}$ curve are more likely caused by imperfect noise modeling for individual pulsars.